PJ is a 50 yo woman admitted to the ICU for acute respiratory failure following abdominal resection for colon cancer. She is 64 inches tall, and weighs 115 lbs. Prior to admission, she had lost 10 pounds over the past month. Her oral intake prior to admission was poor for at least two weeks per her caregivers. Physical exam reveals moderate muscle wasting to her temples and clavicles, and moderate bilateral edema to her legs. She is currently NPO due to ventilator status.   

Using the AND/ASPEN consensus statement, does she meet criteria for diagnosis of malnutrition? If so, what degree of malnutrition would you rank her?

Assuming an NG tube was placed for nutrition support, determine the patient’s energy and protein needs and write a TF order for this patient.
Answer #1

1a) Yes ,she meets the two criteria for diagnosis of malnutrition are:-

- severe muscle wasting

- bilateral pitting edema

1b) Women's age is 50years

Height is 64inches equal to 5feet 4inches

The ideal weight of women of 50year is 100pound with a height of 5feet ,and plus 5pounds for each inches of height over 5feet

So, the weight for height should be 120 pounds but the weight is 115pounds only

According to Gomez classification , all cases of edema are considered as 3rd degree irrespective of weight for age .

Weight for age %= 3rd degree means less than 60

1c) Total protein requirement for adults a day is 0.75gm per kg body weight = 0.75×( 115pound × 0.454)

= 0.75× 52 = 39gm protein is required by the patient .

Total energy requirement is calculated by :-

First calculate the resting metabolic rate for women

= 665+(9.56×weight in kg) + (1.85×height in cms ) -(4.68×age in years)

Weight is 115pounds= 52kg

Height is 5feet 4inches = 162.5cms

Age is 50 years

= 665+(9.56× 52)+ ( 1.85× 162.5) - ( 4.68× 50)

= 665+ ( 497)+ ( 300.6)- (234)

= 1228.6kcal/day × physical activity

Since the patient is on ventilator her physical activity is sedentary so we take 1.4

= Daily energy requirement = 1228.6×1.4 = 1720kcal/day

1d) Patient needs 1720 kcal / day and considering that the patient is receiving a standard formula that provides 1.5kcal/ml . Then the total volume of formula required is = 1720/1.5 = 1146.6ml

If the patient need to receive intermittent feed 6 times a day then ,1146.6ml/ 6 = 191.1ml of formula at each feed.